I am not sure how to solve this:
2d^2 + 3d^2 - 4e^2 - 5e^2
I know (2d^2 + 3d^2) = 5d^2
but is the next part of the equation
-(4e^2 - 5e^2) = -1e^2 or
+(-4e^2 - 5e^2) = -9e^2
Can you please help?
Thank you!!!!

Found 2 solutions by stanbon, rapaljer:
Answer by stanbon(75887)

2d^2 + 3d^2 - 4e^2 - 5e^2
I know (2d^2 + 3d^2) = 5d^2
but is the next part of the equation
-(4e^2 - 5e^2) = -1e^2 or
+(-4e^2 - 5e^2) = -9e^2
-------------------------------
-4e^2 -5e^2 = -9e^2
======================================
Cheers,
Stan H.

Answer by rapaljer(4671)

- 4e^2 - 5e^2 would be -9e^2

R^2